OpenStudy (anonymous):

find g'(x) using the limit method where g(x)=sqrt(1-3x)

OpenStudy (anonymous):

not sure what limit method means... i can only think of using the definition of a derivative

OpenStudy (anonymous):

\[\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

it's just finding the limit of g(x)=sqrt(1-3x)

OpenStudy (anonymous):

yeahh that one. but how do i do it?

OpenStudy (anonymous):

\[\lim_{h\rightarrow 0}\frac{\sqrt{1-3(x+h)}-\sqrt{1-3x}}{h}\]

OpenStudy (anonymous):

\[\lim_{h\rightarrow 0}\frac{\sqrt{1-3x+3h}-\sqrt{1-3x}}{h}\]

OpenStudy (anonymous):

okay i understand that...but what happens next?

OpenStudy (anonymous):

will the root always be there until the end?

OpenStudy (anonymous):

no you need to get it out of the square root

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

\[\lim_{h\rightarrow 0}\frac{\sqrt{1-3x+3h}-\sqrt{1-3x}*(\sqrt{1-3x+3h}+\sqrt{1-3x})}{h(\sqrt{1-3x+h}+\sqrt{1-3x})}\] \[\frac{1-3x+3h-1+3x}{h(\sqrt{1-3x+3h}+\sqrt{1-3x})}\] \[\frac{3h}{h\sqrt{1-3x+3h}+\sqrt{1-3x}}\] \[\frac{3}{\sqrt{1-3x+3h}+\sqrt{1-3x}}\] taking the limit you get \[\frac{3}{2\sqrt{1-3x}}\]

OpenStudy (anonymous):

which is correct because if you use chain rule if you know it \[y'=\frac{1}{2}((\sqrt{1-3x})^{-1/2}*(3)\] which is the same as above

OpenStudy (anonymous):

all i did was multiply by the conjugate

OpenStudy (anonymous):

ohhh. i remember that now. thank you so much for doing all of this. it was really helpful. =)

OpenStudy (anonymous):

no problem =]